If an old guy does a job in 60 hours and an old guy and a new guy does a job in 15 hours, how long does the new guy do it himself?
Answer by stanbon(75887)

If an old guy does a job in 60 hours and an old guy and a new guy does a job in 15 hours, how long does the new guy do it himself?
-------
old guy rate= 1/60 job/hr
new guy rate= 1/x job/hr
Together rate= 1/15 job/hr
------------
Equation:
rate + rate = together rate
1/60 + 1/x = 1/15
x + 60 = 4x
3x = 60
x = 20 hrs (time for the new guy to do the job alone)
==============================
Cheers,
Stan H.
=============